# Learning in a Multifield Coherent Ising Machine

### v2

@@ Abstract
We introduce a network of coupled oscillators that can learn to solve a classification task from a set of examples—performing both training and inference through the nonlinear evolution of the system. We accomplish this by combining three key elements to achieve learning: A long-term memory that stores learned responses, analogous to the synapses in biological brains; a short-term memory that stores the neural activations, similar to the firing patterns of neurons; and an evolution law that updates the synapses in response to novel examples, inspired by synaptic plasticity. Achieving all three elements in wave-based information processors such as metamaterials is a significant challenge. Here, we solve it by leveraging the material multistability to implement long-term memory, and harnessing symmetries and thermal noise to realize the learning rule. Our analysis reveals that the learning mechanism, although inspired by synaptic plasticity, also shares parallelisms with bacterial evolution strategies, where mutation rates increase in the presence of noxious stimuli.
@@ Introduction
Figure 1: Self-learning metamaterial (a) During inference, the metamaterial acts as a mass-spring network. Features are encoded in the excitation amplitude at designated input sites (yellow and purple spheres). If the output amplitude exceeds a threshold, an object is recognized. (b) For the material to learn according to a contrastive learning rule, an identical copy to the lattice in (a) is constructed. The original lattice is operated normally, letting the output site vibrate freely, and is referred to as the free lattice. In the copy, the output is clamped to vibrate at the correct amplitude. At every learning iteration, the coupling springs are updated depending on the difference between the vibrating amplitude of the clamped and free lattices. (c) In the proposed metamaterial, the masses of the free and clamped lattices are mapped to normal modes of each metamaterial site. We refer to these as $\psi_{x}$ and $\psi_{y}$ respectively (collectivelly referred as computational degrees of freedom). The coupling springs are encoded in an additional mode, that is represented by $\sigma$. This mode is parametrically driven at twice its natural frequency (dotted green line) and consequently has two stable phases of oscillation (represented as spin states with blue and red arrows). As a consequence of a nonlinear interaction potential, the effective springs connecting the computational modes depend on the weight state—taking a weak value when neighboring strings are aligned, and a strong value when neighboring weights are oscillating in opposite phases. The learning potential (blue line connecting the modes) causes the weight state to change when clamped and free copies disagree, implementing a contrastive learning rule.
An open problem in physical learning is figuring out how to build physical systems where the learning occurs through the natural evolution of the material, instead of relying on manual updates or external electronic circuitry. Such material would extend the parallelism and speed of wave computing to the learning process, enabling the use of physical phenomena to train machine learning models. However, realizing self-learning in metamaterials is a significant challenge. This is because a self-learning material is tasked with solving two simultaneous problems [24]: On one hand, the material performs a direct problem on the input signals. This direct problem can be, for example, solving a classification task. On the other hand, the material is also performing an inverse problem, autonomously adapting its parameters to improve its fitness in response to new examples.
@@ Results
The requirement for two copies of the system has been an obstacle to the realization of contrastive learning schemes [29]. Here, we overcome this challenge by relying on the crystalline symmetries of the metamaterial (Fig. 1c). When a lattice exhibits a symmetry, such as the four-fold rotational symmetry $C_{4}$ present in our case, eigenmodes corresponding to two-dimensional irreducible representations of the symmetry automatically appear in pairs [30]. Around the frequency of these eigenmodes, the system effectively behaves as two separate, identical oscillator networks. These two networks can be addressed independently by controlling the location at which excitation and clamping are applied. We will use these two symmetry-protected, identical copies to construct the free and clamped copies required by the contrastive learning rule. We will denote the modal displacements at site $i$ as by ${\psi_{i}}_{x}$ and ${\psi_{i}}_{y}$ respectively, and refer to both of them together as the computational field $\vec{\psi}$.
@@
We set the frequency of the computational and learning degrees of freedom to $\omega_{c}^{2}=0.5$ and $\omega_{l}^{2}=1$ respectively. The corresponding quality factors are set to $Q_{c}=4000$ and $Q_{l}=40$, while the parametric driving strength—how much the resonator frequency is externally varied over time—is set to $\alpha_{l}=0.1$. The nonlinear Duffing coefficient plays a critical role in preventing the amplitude of the mode $\sigma_{i}$ to grow without bounds, and is set to $\epsilon=0.005$. The terms $\xi$ represent the thermal noise exerted by the environment on the system. They take the usual form of a set of independent, Gaussian distributed random forces with autocorrelation $\langle\xi_{l/c}(t_{0})\xi_{l/c}(t_{1})\rangle=\sqrt{2k_{B}Tb_{c/l}}\delta(t_{1}-t_{0})$. Interactions between degrees of freedom are captured by the learning potential $H_{l}=H_{I}+H_{L}$, where the local part $H_{L}$ couples the fields locally at every site of the lattice—updating the long-term memory according to the learning rule—and the interaction part $H_{I}$ couples each site with its nearest neighbors—ensuring that the field $\vec{\psi}$ processes information according to the weights stored in the $\sigma$ field.
Figure 2: Single-site learning dynamics. (a) Evolution of the energies in the weight ($\sigma$, middle panel) and computational ($\psi_{x}$, $\psi_{y}$, bottom panel) resonators, as a learning pulse is applied (top panel). b Probability of flip during a learning protocol as a function of $\psi_{X}-\psi_{Y}$ for different temperatures, corresponding to $k_{B}T$ values of $0.001$ (blue), $0.05$ (orange) and $0.1$ (green). c Parametric self-oscillation region (shaded grey area) of the weight degree of freedom $\sigma$; in the shaded region, the weight is bistable and thus has long term memory. The excitation conditions are shown as dashed orange lines and red dot. When $\psi_{X}-\psi_{Y}$ is not zero, applying the learning potential shifts the self-oscillation region. The value of $\psi_{x}-\psi_{y}$ where $\sigma$ goes out of parametric resonance is shown as a purple dashed line in panel b. d Probability of spin flip as a function of $\psi_{x}-\psi_{y}$, for learning potentials $\mu_{m}=\mu_{m,0}/n^{2}$ with $n=1$ (blue), $n=2$ (orange) and $n=3$ (green), illustrating how the step location can be shifted by setting the potential. e Probability of spin flip as a function of the protocol durations $\Delta_{t}=50$ (blue), $\Delta_{t}=100$ (orange), and $\Delta_{t}=300$ (green), showing the emergence of probability oscillations at short learning pulses. The dashed curves in panels c and d show the fit with Eq. 5. Throughout the figure, the parameters are $2\omega_{l}=2$, $\omega_{l}^{2}\alpha=0.1$, $\mu_{m,0}=0.005$, and the step duration is $\Delta_{T}=300$ unless stated otherwise.
@@
The difference between clamped and free oscillations affects the probability of flipping a weight in a step-like fashion (Fig. 2b). This step-like response arises because $\sigma_{i}$ only self-oscillates in a specific region of parametric excitation frequencies and amplitudes (Fig. 2c). When the difference between clamped and free oscillators is large enough, the nonlinear interaction (Eq. 3) shifts this region outside of the excitation conditions. As a consequence, the system ceases to self-oscillate and forgets its state. At low temperature, this phenomenon is abrupt—as long as the excitation remains within the self-oscillation region, the system remembers its state. In contrast, when the temperature is finite, random thermal flips start to occur at finite amplitudes, smoothing out the response (Fig. 2b). The difference $\psi_{x}-\psi_{y}$ at which the forgetting transition occurs is a tunable parameter. We can control it by adjusting the strength of the interaction $\mu_{m}$ (Fig. 2d).
Figure 3: Site-site interactions in a lattice (a) Example metamaterial, consisting of a 4x4 lattice. A harmonic force with frequency $\omega_{c}$ and amplitude $F_{i}$ is applied to both clamped and free degrees of freedom at the input site, situated at the top-left corner. The output amplitude $A_{o}$ is measured at the bottom-right corner. (b) Cumulative density of states (orange) and density of states (blue) as a function of the transmissivity, computed via the effective mass-spring model. The dots correspond to simulations of the transmissivity using the full nonlinear system (Eq. 2) with an excitation force of $10^{-4}$. (c) Weight configurations corresponding to the dots in (b) (d) Evolution of the transmissivity as a function of training iteration for target values $A_{T}$ of 0, 11 and 22 (dashed line). The line represents the mean of N simulations, while the shaded area represents a standard deviation. (e) Output amplitude $A_{o}$ as a function of the clamping amplitude, after 200 learning iterations, starting from a random configuration. The shaded area represents one standard deviation. The dashed orange line corresponds to an ideal learning response. We observe that, for transmisivitty values for which a weight configuration exists, the transmitted amplitude approximately converges to the clamping amplitude.
@@
This term generates the effective coupling springs that connect neighboring computational degrees of freedom ($\psi_{x,i}\leftrightarrow\psi_{x,j}$ and $\psi_{y,i}\leftrightarrow\psi_{y,j}$). Depending on the phase of the weight degrees of freedom, these coupling springs can take two values (Fig. 1)—determined by $\lambda(t)=\lambda_{m}$ and $c_{0}$; if the weight field has the same phase of oscillation in both sites, the spring constant will be $k_{low}=\lambda_{m}c_{0}$, while if weights oscillate at opposing phases, the spring constant will be $k_{high}=\lambda_{m}(c_{0}+4<\sigma^{2}>)$. This effective spring model is valid as long as the computational amplitude remains small and the potential in Eq. 6 does not alter the weight amplitude.
Although each individual weight $\sigma_{i}$ is binary, in metamaterials composed of multiple sites (Fig. 3a) the transmitted energy between input and output can take many different values, depending on the global weight configuration. This can be seen in the density of states, that quantifies the number of weight configurations that result in a lattice transmissivity around a specific value (Fig. 3b,c). Consequently, such lattice can be used to approximate continuous transmissivities, as long as they fall within the range of expressible values.
@@
We illustrate the capabilities of the proposed system on the Iris classification test [33]. The dataset contains measurements of the petal length $l_{p}$, petal width $w_{p}$, sepal length $l_{s}$ and sepal width $s_{w}$ (Fig. 4a) for a set of 150 flowers belonging to the species iris setosa, iris virginica, and iris versicolor. The classifier consists of a system of three 13x13 lattices; the lattice with the highest output amplitude, measured at the central site, will be the inferred flower species. We encode the geometric features ($l_{p}$, $w_{p}$, $l_{s}$, $w_{s}$) in the amplitudes of a set harmonic excitation forces acting on each lattice. These signals are applied at sites through the boundary of the lattice (Fig. 4b), acting on both free $\psi_{x}$ and clamped $\psi_{y}$ computational modes. Since the lattices cannot produce large negative transmissivities (Fig. 3b), we apply both positive and negative copies of each feature signal.
@@
The training process starts by initializing each of the three lattices with a random weight configuration. Then, for every learning iteration, we subject all lattices to the forces encoding the features of a randomly chosen flower. We clamp the output site $\psi_{y,center}$ to $55$ for the lattice corresponding to the current type of flower, and to 35-35 for the lattices corresponding to the other flowers. Finally, we apply a weight update protocol to all lattices, consisting of setting the coupling $\lambda$ to zero, and changing $\mu(t)$ following a Gaussian pulse. This process is repeated for 2000 iterations.
We observe that the system starts from an accuracy of $33.2\pm 22.1\,\%$ , as one would expect from random choice. As more training iterations are performed, the accuracy increases to $71.9\pm 9.8\,\%$, providing a clear signature of physical learning. These results are remarkable as the system stores only 156 bits of information, contains only Kerr nonlinearities, and is directly trained by exposing it to the desired output. Using a more advanced training algorithm, where the learning process is stopped when the precision exceeds a threshold, we achieve $80.0\pm 7.0\,\%$ accuracy, with the best-performing runs reaching $87.3\,\%$. We hypothesize that larger, architected networks trained under more complex protocols, such those where the target value for every sample is adjusted according to the lattice output [23], the system could be able to reach performances comparable with electronic linear classifiers.

### v3

# Multimodal oscillator networks learn to solve a classification problem

@@ Abstract
We numerically demonstrate a network of coupled oscillators that can learn to solve a classification task from a set of examples—performing both training and inference through the nonlinear evolution of the system. We accomplish this by combining three key elements to achieve learning: A long-term memory that stores learned responses, analogous to the synapses in biological brains; a short-term memory that stores the neural activations, similar to the firing patterns of neurons; and an evolution law that updates the synapses in response to novel examples, inspired by synaptic plasticity. Achieving all three elements in wave-based information processors such as metamaterials is a significant challenge. Here, we solve it by leveraging the material multistability to implement long-term memory, and harnessing symmetries and thermal noise to realize the learning rule. Our analysis reveals that the learning mechanism, although inspired by synaptic plasticity, also shares parallelisms with bacterial evolution strategies, where mutation rates increase in the presence of noxious stimuli.
@@ Introduction
Figure 1: Self-learning metamaterial (a) During inference, the metamaterial acts as a mass-spring network. Features are encoded in the excitation amplitude at designated input sites (yellow and purple spheres). If the output amplitude exceeds a threshold, an object is recognized. (b) For the material to learn according to a contrastive learning rule, an identical copy to the lattice in (a) is constructed. The original lattice is operated normally, letting the output site vibrate freely, and is referred to as the free lattice. In the copy, the output is clamped to vibrate at the correct amplitude. At every learning iteration, the coupling springs are updated depending on the difference between the vibrating amplitude of the clamped and free lattices. (c) In the proposed metamaterial, the masses of the free and clamped lattices are mapped to normal modes of each metamaterial site. We refer to these as $\psi_{x}$ and $\psi_{y}$ respectively (collectively referred as computational degrees of freedom). The coupling springs are encoded in an additional mode, that is represented by $\sigma$. This mode is parametrically driven at twice its natural frequency (dotted green line) and consequently has two stable phases of oscillation (represented as spin states with blue and red arrows). As a consequence of a nonlinear interaction potential, the effective springs connecting the computational modes depend on the weight state—taking a weak value when neighboring strings are aligned, and a strong value when neighboring weights are oscillating in opposite phases. The learning potential (blue line connecting the modes) causes the weight state to change when clamped and free copies disagree, implementing a contrastive learning rule.
An open problem in physical learning is figuring out how to build physical systems where the learning occurs through the natural evolution of the material, instead of relying on manual updates or external electronic circuitry. Such material could potentially extend the parallelism and speed of wave computing to the learning process, enabling the use of physical phenomena to train machine learning models. However, realizing self-learning in metamaterials is a significant challenge. This is because a self-learning material is tasked with solving two simultaneous problems [24]: On one hand, the material performs a direct problem on the input signals. This direct problem can be, for example, solving a classification task. On the other hand, the material is also performing an inverse problem, autonomously adapting its parameters to improve its fitness in response to new examples.
@@ Results
The requirement for two copies of the system has been an obstacle to the realization of contrastive learning schemes [29]. Here, we overcome this challenge by relying on the crystalline symmetries of the metamaterial (Fig. 1c). When a lattice exhibits a symmetry, such as the four-fold rotational symmetry $C_{4}$ present in our case, eigenmodes corresponding to two-dimensional irreducible representations of the symmetry automatically appear in pairs [30]. Around the frequency of these eigenmodes, the system effectively behaves as two separate, identical oscillator networks. These two networks can be addressed independently by controlling the location at which excitation and clamping are applied. We will use these two symmetry-protected, identical copies to construct the free and clamped copies required by the contrastive learning rule. We will denote the modal displacements at site $i$ by ${\psi_{i}}_{x}$ and ${\psi_{i}}_{y}$ respectively, and refer to both of them together as the computational field $\vec{\psi}$.
@@
We set the frequency of the computational and learning degrees of freedom to $\omega_{c}^{2}=0.5$ and $\omega_{l}^{2}=1$ respectively (the subindex is used to distinguish between computational and learning modes). The corresponding quality factors are set to $Q_{c}=4000$ and $Q_{l}=40$, while the parametric driving strength—how much the resonator frequency is externally varied over time—is set to $\alpha_{l}=0.1$. The nonlinear Duffing coefficient plays a critical role in preventing the amplitude of the mode $\sigma_{i}$ to grow without bounds, and is set to $\epsilon=0.005$. The terms $\xi$ represent the thermal noise exerted by the environment on the system. They take the usual form of a set of independent, Gaussian distributed random forces with autocorrelation $\langle\xi_{l/c}(t_{0})\xi_{l/c}(t_{1})\rangle=\sqrt{2k_{B}Tb_{c/l}}\delta(t_{1}-t_{0})$. Interactions between degrees of freedom are captured by the learning potential $H_{l}=H_{I}+H_{L}$, where the local part $H_{L}$ couples the fields locally at every site of the lattice—updating the long-term memory according to the learning rule—and the interaction part $H_{I}$ couples each site with its nearest neighbors—ensuring that the field $\vec{\psi}$ processes information according to the weights stored in the $\sigma$ field.
Figure 2: Single-site learning dynamics. (a) Evolution of the energies in the weight ($\sigma$, middle panel) and computational ($\psi_{x}$, $\psi_{y}$, bottom panel) resonators, as a learning pulse is applied (top panel). (b) Probability of flip during a learning protocol as a function of $\psi_{X}-\psi_{Y}$ for different temperatures, corresponding to $k_{B}T$ values of $0.001$ (blue), $0.05$ (orange) and $0.1$ (green). (c) Parametric self-oscillation region (shaded grey area) of the weight degree of freedom $\sigma$; in the shaded region, the weight is bistable and thus has long term memory. The excitation conditions are shown as dashed orange lines and red dot. When $\psi_{X}-\psi_{Y}$ is not zero, applying the learning potential shifts the self-oscillation region. The value of $\psi_{x}-\psi_{y}$ where $\sigma$ goes out of parametric resonance is shown as a red dashed line in panel b. (d) Probability of spin flip as a function of $\psi_{x}-\psi_{y}$, for learning potentials $\mu_{m}=\mu_{m,0}/n^{2}$ with $n=1$ (blue), $n=2$ (orange) and $n=3$ (green), illustrating how the step location can be shifted by setting the potential. (e) Probability of spin flip as a function of the protocol durations $\Delta_{t}=50$ (blue), $\Delta_{t}=100$ (orange), and $\Delta_{t}=300$ (green), showing the emergence of probability oscillations at short learning pulses. The dashed curves in panels (c) and (d) show the fit with Eq. 5. Throughout the figure, the parameters are $2\omega_{l}=2$, $\omega_{l}^{2}\alpha=0.1$, $\mu_{m,0}=0.005$, and the step duration is $\Delta_{T}=300$ unless stated otherwise.
@@
Although here we chose to apply the learning potential $\mu(t)$ according to a Gaussian pulse, its specific shape is not critical. In the Appendix, we show that following a square pulse results in similar dynamics. Because $\mu(t)$ is independent of the state of the system, it is the same for every resonator and does not require a controller; its role is that of a system clock, indicating when learning shall occur.
@@
The difference between clamped and free oscillations affects the probability of flipping a weight in a step-like fashion (Fig. 2b). This step-like response arises because $\sigma_{i}$ only self-oscillates in a specific region of parametric excitation frequencies and amplitudes (Fig. 2c). When the difference between clamped and free oscillators is large enough, the nonlinear interaction (Eq. 3) shifts this region outside of the excitation conditions. As a consequence, the system ceases to self-oscillate and forgets its state. At low temperature, this phenomenon is abrupt—as long as the excitation remains within the self-oscillation region, the system remembers its state. In contrast, when the temperature is finite, random thermal flips start to occur at finite amplitudes, smoothing out the response (Fig. 2b). The difference $\psi_{x}-\psi_{y}$ at which the forgetting transition occurs is a tunable parameter. We can control it by adjusting the strength of the interaction $\mu_{m}$ (Fig. 2d). This state-dependent switching relies on the ratio between the quality factors of the computational and learning degrees of freedom. Because outside parametric resonance the computational degree of freedom has a much longer memory lifetime than the weight degree of freedom, the difference between clamped and free copies remains throughout the update protocol. In a practical scenario, this difference could be implemented by controlling the coupling between each mode and the environment, or via external cooling signals to speed up the learning dynamics.
Figure 3: Site-site interactions in a lattice (a) Example metamaterial, consisting of a 4x4 lattice. A harmonic force with frequency $\omega_{c}$ and amplitude $F_{i}$ is applied to both clamped and free degrees of freedom at the input site, situated at the top-left corner. The output amplitude $A_{o}$ is measured at the bottom-right corner. (b) Cumulative density of states (orange) and density of states (blue) as a function of the transmissivity, computed via the effective mass-spring model. The dots correspond to simulations of the transmissivity using the full nonlinear system (Eq. 2) with an excitation force of $10^{-4}$. (c) Weight configurations corresponding to the dots in (b) (d) Evolution of the transmissivity as a function of training iteration for target values $A_{T}$ of 0, 11 and 22 (dashed line). (e) Output amplitude $A_{o}$ as a function of the clamping amplitude, after 200 learning iterations, starting from a random configuration. The shaded area represents one standard deviation. The dashed orange line corresponds to an ideal learning response. We observe that, for transmissivity values for which a weight configuration exists, the transmitted amplitude approximately converges to the clamping amplitude. Panels (d) and (e) have been computed by averaging 2000 training runs, with the shaded area representing the standard deviation.
@@
This term generates the effective coupling springs that connect neighboring computational degrees of freedom ($\psi_{x,i}\leftrightarrow\psi_{x,j}$ and $\psi_{y,i}\leftrightarrow\psi_{y,j}$). Depending on the phase of the weight degrees of freedom, these coupling springs can take two values (Fig. 1)—determined by $\lambda(t)=\lambda_{m}$ and $c_{0}$; if the weight field has the same phase of oscillation in both sites, the spring constant will be $k_{low}=\lambda_{m}c_{0}$, while if weights oscillate at opposing phases, the spring constant will be $k_{high}=\lambda_{m}(c_{0}+4<\sigma^{2}>)$. This effective spring model is valid as long as the computational amplitude remains small and the potential in Eq. 6 does not alter the weight amplitude. Similarly to $\mu(t)$, $\lambda(t)$ is a state-independent signal that is identical for all resonators, acting as a system clock.
Although each individual weight $\sigma_{i}$ is binary, in metamaterials composed of multiple sites (Fig. 3a) the transmitted energy between input and output can take many different values, depending on the global weight configuration. This can be seen in the density of states, which quantifies the number of weight configurations that result in a lattice transmissivity around a specific value (Fig. 3b,c). This distribution depends on lattice size and on the $k_{high}/k_{low}$ ratio, which is chosen (approximately, by trial and error) to have the transmissivities spread out as uniformly as possible. Because the lattice can be trained to approximate continuous transmissivities, as long as they fall within the range of expressible values, we treat this element as a single programmable weight—which we refer to as a learnistor. Networks of trainable weights, combined with fixed nonlinear activation functions, are computationally universal [33]; thus, we hypothesize that such trainable elements can form building blocks for more advanced learning architectures.
@@
We illustrate the capabilities of the proposed system on the Iris classification test [34]. The dataset contains measurements of the petal length $l_{p}$, petal width $w_{p}$, sepal length $l_{s}$ and sepal width $w_{s}$ (Fig. 4a) for a set of 150 flowers belonging to the species iris setosa, iris virginica, and iris versicolor. The classifier consists of a system of three 13x13 lattices; the lattice with the highest output amplitude, measured at the central site, will be the inferred flower species. We encode the geometric features ($l_{p}$, $w_{p}$, $l_{s}$, $w_{s}$) in the amplitudes of a set of harmonic excitation forces acting on each lattice. These signals are applied at sites through the boundary of the lattice (Fig. 4b), acting on both free $\psi_{x}$ and clamped $\psi_{y}$ computational modes. The input and output locations are selected so that the output is at an approximately equal distance from all inputs. Since the lattices cannot produce large negative transmissivities (Fig. 3b), we apply both positive and negative copies of each feature signal. Alternative interaction potentials can actually produce negative transmissivities, as is discussed in the appendix.
@@
The training process starts by initializing each of the three lattices with a random weight configuration. Then, for every learning iteration, we subject all lattices to the forces encoding the features of a randomly chosen flower. We clamp the output site $\psi_{y,center}$ to $55$ for the lattice corresponding to the current type of flower, and to 35-35 for the lattices corresponding to the other flowers (these values have been chosen empirically to fall within the range of expressible values). Finally, we apply a weight update protocol to all lattices, consisting of setting the coupling $\lambda$ to zero, and changing $\mu(t)$ following a Gaussian pulse. This process is repeated for 2000 iterations.
We observe that the system starts from an accuracy of $33.2\pm 22.1\,\%$ , as one would expect from random choice. As more training iterations are performed, the accuracy increases to $71.9\pm 9.8\,\%$, providing a clear signature of physical learning. These results are remarkable as the system stores only 156 bits of information, contains only Kerr nonlinearities, and is directly trained by exposing it to the desired output. Using a more advanced training algorithm, where the learning process is stopped when the precision exceeds a threshold, we achieve $80.0\pm 7.0\,\%$ accuracy, with the best-performing runs reaching $87.3\,\%$. Although these accuracy figures are much lower than traditional digital machine learning algorithms (methods such as k-nearest neighbors, support vector machines and decision trees achieve performances in the $93.0-97\,\%$ range), and to electronic implementations of learning rules (which can already reach values around $95\,\%$), they are remarkable in that they are achieved through the dynamical evolution of a nonlinear mass-spring network (subject to examples and simple clocking pulses). Our accuracies are however in line with similar ’unconventional learning’ demonstrations, such as classifying with origami sheets (which reach best-values of $92\,\%$ on a simpler two-class flower test) or when using EEG signals (which achieve accuracies $80\,\%$). To confirm that our classification provides evidence of learning, and is not a result of random fluctuations, we perform statistical hypothesis testing in the Appendix.
Although in line with other unconventional learning results, the limited accuracy of our trained classifier merits further investigation. We identify two possible limiting factors: The binary nature of the weights in our system—which may reduce the model expressivity, or peculiarities of the training protocol. To separate the two effects, we construct a digitally-trained multiclass linear architecture using networks of $3x3$ resonator lattices, which are highly quantized. The model projects a feature vector $\vec{f}=(p_{l},p_{w},1)$ onto a set of weight vectors corresponding to the three flower classes $\vec{w}_{virginica},\vec{w}_{versicolor},\vec{w}_{setosa}$}. Every class is implemented using a network of three pairs of $3x3$ lattices connected to an output mass (Fig. 4e), with each pair of lattices expressing the positive and negative values of a specific weight vector element. Using such architecture, as opposed to a large lattice, allows us to identify specific sub-lattices with model coefficients. We initialize the spin texture of every lattice to the closest value required by the linear model. When simulating the lattice dynamics, the model achieves an accuracy of $97.3\,\%$—the maximum possible for a linear classifier. This indicates that weight quantization cannot explain the observed performance degradation
Therefore, the low precision on the Iris dataset must originate in the training dynamics. We know from Fig. LABEL:fig:b, that lattices can be trained to arbitrary transmissivities within their range of expressivity. This works because spin changes stop taking place when the transmitted amplitude matches the target. However, when training on the Iris dataset, each sample causes the lattice to shift towards a different transmissivity (Fig. 4f), preventing the lattice from settling into a stable value. In prior, high-accuracy, electronic realizations of contrastive learning [23], the training process has not prescribed the output to a specific value, but instead nudged incorrect answers towards the correct value. We hypothesize that such sophisticated training methods may enable the system to achieve higher accuracies—potentially comparable to electronic linear classifiers, by allowing it to settle into high-accuracy configurations even if the output does not match a specific transmitted amplitude. However, evaluating such more complex protocols falls outside the scope of this work, which focuses on control-free systems, i.e. where the input signals applied to the network are independent of its state.